# Self-Supervised Depth Estimation Based on Camera Models

### v1

@@ Abstract
Depth estimationn is a critical topic for robotics and vision-related tasks. In monocular depth estimation, in comparison with supervised learning that requires expensive ground truth labeling, self-supervised methods possess great potential due to no labeling cost. However, self-supervised learning still has a large gap with supervised learning in depth estimation performance. Meanwhile, scaling is also a major issue for monocular unsupervised depth estimation, which commonly still needs ground truth scale from GPS, LiDAR, or existing maps to correct. In deep learning era, while existing methods mainly rely on the exploration of image relationships to train the unsupervised neural networks, fundamental information provided by the camera itself has been generally ignored, which can provide extensive supervision information for free, without the need for any extra equipment to provide supervision signals. Utilizing the camera itself’s intrinsics and extrinsics, depth information can be calculated for ground regions and regions connecting ground based on physical principles, providing free supervision information without any other sensors. The method is easy to realize and can be a component to enhance the effects of all the unsupervised methods.
@@ I INTRODUCTION
Meanwhile, due to the convenience brought by deep neural networks, extensive information from the sensors themselves has been ignored. This paper introduces a method that leverages camera model parameters (both intrinsic and extrinsic) to accurately calculate depth information, enhancing self-supervised depth estimation. By utilizing the camera parameters, we can precisely determine the depth for much of the scene, facilitating neural network training without needing explicit ground truth data. The method also uses image semantics to compute the ground plane’s depth, allowing for the estimation of objects’ depth on the ground, such as buildings and vehicles. This provides a cost-effective way of depth supervision using only the camera. Our training framework, designed around physics depth, supports and enhances unsupervised networks, allowing for the detailed modeling of 3D structures based on scene depth and camera parameters. Importantly, our algorithm serves as a valuable extension for any unsupervised depth estimation effort.
@@
In summary, our main contributions include the following aspects: 1. We propose a novel mechanism that utilizes camera physics model parameters to calculate depth information for a large portion of the scene, thereby supervising the depth estimation network. We refer to this depth information estimated from the camera physics model as physics depth. 2. To address the uncertain scale issue in unsupervised monocular depth estimation, our approach provides an absolute scale instead of just a relative scale. 3. We designed a neural network training framework to effectively integrate physics depth supervision with unsupervised methods, specifically targeting the physics depth calculated from the camera model. The framework for physics depth computation and training with the self-supervised network is shown in Fig. 1.
@@ III-A Physics depth for Full Field of view
We introduce a novel monocular depth estimation technique called ”physics depth,” which uses camera intrinsic and extrinsic parameters and semantic segmentation to calculate absolute depth. This approach derives depth from basic physical principles, assuming the camera captures initially flat surfaces. We refine depth estimates by identifying truly flat areas through semantic segmentation, extending depth to adjacent areas, and filling gaps with inpainting. Our method, tested on KITTI, CityScape, and Make3D datasets, achieves accuracy comparable to LiDAR, especially for close, flat surfaces. Our model is based on the pinhole camera model, which is prevalent in real-world applications due to its minimal distortion, making it an excellent reference point. Although designed with the pinhole model in mind, our method can be adapted for different camera types by adjusting for each camera’s unique characteristics. For every image pixel, we compute a unit vector $\hat{r}$ representing the camera ray’s direction in the physical world, translating pixel positions into directional vectors that indicate the camera’s viewpoint in real space.
@@ IV-A Network Architecture
Our research addresses the data scarcity in physics depth, which is typically limited to only portions of an image and, by itself, inadequate for self-supervised learning. We enhance depth estimation by using physics depth as an informative prior, enabling our model to refine ground surfaces and surrounding depth estimations more accurately than traditional self-supervised models that initiate with random depth values. This approach significantly improves efficiency in correcting depth inaccuracies through self-supervised training. Our model uniquely combines RGB and physics depth data, adding valuable depth insights. During the supervised phase, we assess the confidence in physics depth to focus learning on the most reliable areas. For self-supervised learning, we advance the model by integrating geometric consistency with photometric consistency, leading to more precise depth estimates.
@@ VI Conclusion
This paper presents a novel self-supervised learning approach that improves monocular depth estimation using physics-based depth cues. Our method surpasses existing self-supervised techniques by incorporating physics depth estimation, enhancing accuracy and environmental detail capture. It allows models to accurately predict scene depth and 3D structures, significantly advancing self-supervised learning with state-of-the-art results on KITTI, Cityscapes, and Make3D datasets. This approach yields better depth information for improved understanding of real-world scenes.

### v2

# Embodiment: Self-Supervised Depth Estimation Based on Camera Models

@@ Abstract
Depth estimationn is a critical topic for robotics and vision-related tasks. In monocular depth estimation, in comparison with supervised learning that requires expensive ground truth labeling, self-supervised methods possess great potential due to no labeling cost. However, self-supervised learning still has a large gap with supervised learning in 3D reconstruction and depth estimation performance. Meanwhile, scaling is also a major issue for monocular unsupervised depth estimation, which commonly still needs ground truth scale from GPS, LiDAR, or existing maps to correct. In the era of deep learning, existing methods primarily rely on exploring image relationships to train unsupervised neural networks, while the physical properties of the camera itself—such as intrinsics and extrinsics—are often overlooked. These physical properties are not just mathematical parameters; they are embodiments of the camera’s interaction with the physical world. By embedding these physical properties into the deep learning model, we can calculate depth priors for ground regions and regions connected to the ground based on physical principles, providing free supervision signals without the need for additional sensors. This approach is not only easy to implement but also enhances the effects of all unsupervised methods by embedding the camera’s physical properties into the model, thereby achieving an embodied understanding of the real world.
@@ I INTRODUCTION
Meanwhile, due to the convenience brought by deep neural networks, extensive information from the sensors themselves has been ignored. This paper introduces a method that leverages camera model parameters (both intrinsic and extrinsic) to accurately calculate depth information, thereby embedding the camera model and its physical characteristics into the deep learning model. This approach goes beyond mere mathematical computation; it embodies the interaction between the camera and the physical world. By integrating these physical characteristics into the deep learning model, we can accurately determine the depth for much of the scene, facilitating neural network training without the need for explicit ground truth data. The method also incorporates image semantics to calculate the ground plane’s depth, allowing for the estimation of the depth of objects on the ground, such as buildings and vehicles. Utilizing this physics-based supervision approach, the framework not only enhances the performance of unsupervised networks but also achieves an embodied understanding of the real world, providing strong support for detailed 3D structure modeling. Importantly, our algorithm serves as a valuable extension for any unsupervised depth estimation effort.
@@
In summary, our main contributions include the following aspects: 1. We propose a novel mechanism that leverages the physical model parameters of the camera to calculate the depth information for a large portion of the scene, embedding the camera model and its physical properties into the deep learning model to supervise the depth estimation network. We refer to this depth information, derived from the camera’s physical model, as physics depth. 2. To address the uncertain scale issue in unsupervised monocular depth estimation, our approach provides an absolute scale instead of just a relative scale. 3. We designed a neural network training framework to effectively integrate physics depth supervision with unsupervised methods, specifically targeting the physics depth calculated from the camera model. The framework for physics depth computation and training with the self-supervised network is shown in Fig. 1.
@@ III-A Physics depth for Full Field of view
We introduce a novel monocular depth estimation technique called ”physics depth”. This method leverages the camera’s intrinsic and extrinsic parameters, as well as semantic segmentation, to calculate absolute depth, embedding the physical characteristics of the camera model into the deep learning model. This approach derives depth from basic physical principles, assuming the camera captures initially flat surfaces. We refine depth estimates by identifying truly flat areas through semantic segmentation, extending depth to adjacent areas, and filling gaps with inpainting. Our method, tested on KITTI, CityScape, and Make3D datasets, achieves accuracy comparable to LiDAR, especially for close, flat surfaces. Our model is based on the pinhole camera model, which is widely used in practical applications due to its minimal distortion, making it an ideal reference point for achieving embodiment. Although designed with the pinhole model in mind, our method can be adapted for different camera types by adjusting for each camera’s unique characteristics. For every image pixel, we compute a unit vector $\hat{r}$ representing the camera ray’s direction in the physical world, translating pixel positions into directional vectors that indicate the camera’s viewpoint in real space.
@@ IV-A Network Architecture
Our research addresses the data scarcity in physics depth, which is typically limited to only portions of an image and, by itself, inadequate for self-supervised learning. By embedding the physical characteristics of the camera model into the deep learning model, we utilize physics depth as an embodied prior, enhancing depth estimation. Unlike traditional self-supervised models that typically start with random depth values, our model can more accurately refine the depth estimation of ground surfaces and surrounding areas. This approach significantly improves efficiency in correcting depth inaccuracies through self-supervised training. Our model uniquely combines RGB and physics depth, adding valuable depth insights. During the supervised phase, we assess the confidence in physics depth to focus learning on the reliable areas. For self-supervised learning, we advance the model by integrating geometric consistency with photometric consistency, leading to more precise depth estimates.
@@ VI Conclusion
This paper presents a novel self-supervised learning approach that integrates the physical characteristics of the camera model with the concept of embodiment, embedding them into the deep learning model. By leveraging physics-based depth cues, our method improves monocular depth estimation. Through the incorporation of physics depth estimation, our approach achieves an embodied understanding of the interaction between the camera and the physical world, enhancing the model’s accuracy and its ability to capture environmental details, surpassing existing self-supervised techniques. This method enables the model to accurately predict scene depth, achieving state-of-the-art self-supervised learning results on the KITTI, Cityscapes, and Make3D datasets. This approach provides better depth information, thereby deepening the model’s embodied understanding of real-world scenes.